# GPT4Scene: Understand 3D Scenes from Videos with Vision-Language Models

### v1

@@ 5.3 Ablation Study
Num Frames Resolution ScanQA ScanRefer
ROUGE CIDEr Acc0.25 Acc0.5
@@ Appendix A Prompts of Closed-source VLMs
Appendix B 2D Multi-modal Benchmark
We tested the fine-tuned model on 2D image and video multimodal large models. Table 7 shows the results of MVBench. As we can see, our model shows improvement for the object and action metrics, indicating that the model fine-tuned with ScanAlign is better at handling spatial variations and the information of objects in the scene. Table 8 presents the results on other benchmarks, where we can observe that after our fine-tuning, the model’s ability to understand images and videos did not decline significantly. It demonstrates the effectiveness of ScanAlign.
Appendix C Qualitative Results
Appendix D Full Quantitive Results
@@
InternVL2-8B [23] 16.9 20.0 9.8 5.2 2.7 32.6 14.5 55.3
MiniCPM-V-2.6 [105] 20.6 23.9 12.0 6.3 8.4 31.5 11.8 60.1
Qwen2-VL-7B [90] 19.0 27.8 13.6 6.3 3.0 34.2 11.4 53.9
Qwen2-VL-7B (GPT4Scene) [90] 25.5 43.4 29.3 20.9 14.6 43.6 17.7 90.9
Qwen2-VL-7B (GPT4Scene)-HD [90] 26.3 41.9 28.6 21.6 15.9 43.6 17.6 89.9
Qwen2-VL-7B (GPT4Scene)-HDM [90] 28.2 44.4 30.3 22.3 15.5 46.5 18.9 96.3
@@
      Qwen2-VL-7B (GPT4Scene) [90]       50.7       70.9       48.0       70.5       52.9       59.3       57.4       60.7
      Qwen2-VL-7B (GPT4Scene)-HD [90]       51.4       69.1       50.2       69.4       51.3       57.9       57.2       60.4
      Qwen2-VL-7B (GPT4Scene)-HDM [90]       55.9       69.9       50.8       68.7       53.3       60.4       59.4       62.4
@@
Qwen2-VL-7B (GPT4Scene) [90] 63.8 36.3 26.5 57.6 60.6 34.2 25.6 55.2
Qwen2-VL-7B (GPT4Scene)-HD [90] 79.1 40.4 28.3 60.2 74.4 37.9 27.3 57.7
Qwen2-VL-7B (GPT4Scene)-HDM [90] 91.7 43.1 29.3 61.9 86.3 40.6 28.2 59.3
@@
Qwen2-VL-7B (GPT4Scene) [90] 65.5 61.2 34.8 31.1 40.5 36.7
Qwen2-VL-7B (GPT4Scene)-HD [90] 77.5 71.9 44.9 40.6 50.9 46.4
Qwen2-VL-7B (GPT4Scene)-HDM [90] 90.3 83.7 56.4 50.9 62.6 57.0
@@
Qwen2-VL-7B (GPT4Scene) [90] 85.2 61.4 60.1 55.1 37.7 34.4 39.4 36.3 45.4 42.1
Qwen2-VL-7B (GPT4Scene)-HD [90] 93.6 81.8 72.5 66.2 46.6 42.9 41.8 38.9 53.7 50.0
Qwen2-VL-7B (GPT4Scene)-HDM [90] 97.4 84.4 85.0 77.7 59.9 55.1 48.6 44.6 64.5 59.8

### v2

@@ 5.3 Ablation Study
Num Resolution ScanQA ScanRefer
Frames ROUGE CIDEr Acc0.25 Acc0.5
@@ Appendix A Prompts of Closed-source VLMs
InternVL2-8B [23] 19.0 23.9 12.0 6.3 3.3 34.3 14.5 62.5
MiniCPM-V-2.6 [105] 20.6 25.1 16.9 12.4 8.4 31.5 11.8 60.1
Qwen2-VL-7B [90] 19.0 27.8 13.6 6.3 3.0 29.3 11.4 53.9
Qwen2-VL-7B (GPT4Scene) 25.5 43.4 29.3 20.9 14.6 43.6 17.7 90.9
Qwen2-VL-7B (GPT4Scene)-HD 26.3 41.9 28.6 21.6 15.9 43.6 17.6 89.9
Qwen2-VL-7B (GPT4Scene)-HDM 28.2 44.4 30.3 22.3 15.5 46.5 18.9 96.3
@@
      Qwen2-VL-7B (GPT4Scene)       50.7       70.9       48.0       70.5       52.9       59.3       57.4       60.7
      Qwen2-VL-7B (GPT4Scene)-HD       51.4       69.1       50.2       69.4       51.3       57.9       57.2       60.4
      Qwen2-VL-7B (GPT4Scene)-HDM       55.9       69.9       50.8       68.7       53.3       60.4       59.4       62.4
@@
Appendix B 2D Multi-modal Benchmark
We tested the fine-tuned Qwen2-VL-7B(GPT4Scene) model on 2D image and video multimodal large models. Table 7 shows the results of MVBench. As we can see, our model shows improvement for the object and action metrics, indicating that the model fine-tuned with ScanAlign is better at handling spatial variations and the information of objects in the scene. Table 8 presents the results on other benchmarks, where we can observe that after our fine-tuning, the model’s ability to understand images and videos did not decline significantly. It demonstrates the effectiveness of ScanAlign.
Appendix C Qualitative Results
Appendix D Full Quantitive Results
@@
Qwen2-VL-7B (GPT4Scene) 63.8 36.3 26.5 57.6 60.6 34.2 25.6 55.2
Qwen2-VL-7B (GPT4Scene)-HD 79.1 40.4 28.3 60.2 74.4 37.9 27.3 57.7
Qwen2-VL-7B (GPT4Scene)-HDM 91.7 43.1 29.3 61.9 86.3 40.6 28.2 59.3
@@
Qwen2-VL-7B (GPT4Scene) 65.5 61.2 34.8 31.1 40.5 36.7
Qwen2-VL-7B (GPT4Scene)-HD 77.5 71.9 44.9 40.6 50.9 46.4
Qwen2-VL-7B (GPT4Scene)-HDM 90.3 83.7 56.4 50.9 62.6 57.0
@@
Qwen2-VL-7B (GPT4Scene) 85.2 61.4 60.1 55.1 37.7 34.4 39.4 36.3 45.4 42.1
Qwen2-VL-7B (GPT4Scene)-HD 93.6 81.8 72.5 66.2 46.6 42.9 41.8 38.9 53.7 50.0
Qwen2-VL-7B (GPT4Scene)-HDM 97.4 84.4 85.0 77.7 59.9 55.1 48.6 44.6 64.5 59.8